OpenStudy (anonymous):

What is the minimum or maximum value of the function y=2x^2-8x+5? Please help me fast!!!

OpenStudy (zzr0ck3r):

4x - 8 = 0, solve and you have the answer

OpenStudy (zzr0ck3r):

this will be the x coordinate, plug it into the original function to get the value

OpenStudy (anonymous):

can you please do it. I really don't have time do it.

OpenStudy (anonymous):

For the Xv=-b/2a and for Yv=-(b^2-ac)/4a

OpenStudy (anonymous):

This is for an Algebra 2 take home test. I'm currently failing. So I really need your help.